OpenStudy (anonymous):

[8.03] Using complete sentences, explain how to completely factor 2x2 − 6x − 36.

OpenStudy (anonymous):

Factor out 2. \[2 x^2-6 x-36=2\left( x^2-3 x-18\right) \]Factor the quadratic.\[2\left( x^2-3 x-18\right)=2 (x-6) (x+3) \]